The sum of three consecutive integers is 53 more than the least of the integers. Find the integers.

Answer by checkley77(12844)

Let x, x+1 & x+2 be the three integers.
x+x+1+x+2=x+53
3x+3=x+53
3x-x=53-3
2x=50
x=50/2
x=25 fir the first integer.
25=1=26 for the next integer.
25+2=27 is the largest integer.
Proof:
25+25+27=25+53
77=77